# Exploring SMEFT Couplings Using the Forward-Backward Asymmetry in Neutral Current Drell-Yan Production at the LHC

### v2

@@ Abstract
Neutral current Drell-Yan (DY) lepton-pair production is considered in the framework of the Standard Model Effective Field Theory (SMEFT). Using the open-source fit platform xFitter, we investigate the impact of high-statistics measurements of the neutral current DY (NCDY) forward-backward asymmetry $A_{\rm{FB}}$ near the weak boson mass scale in the present and forthcoming stages of the Large Hadron Collider (LHC). Besides recovering earlier results on the $A_{\rm{FB}}$ sensitivity to parton distribution functions, we analyze the precision determination of $Z$-boson couplings to left-handed and right-handed $u$-quarks and $d$-quarks, and explore Beyond-Standard-Model contributions using the SMEFT framework. We perform a sensitivity study and comment on the role of the $A_{\rm{FB}}$ asymmetry for the electroweak SMEFT fit and precision $Z$-boson physics at the LHC and high-luminosity HL-LHC.
@@ 1 Introduction
Furthermore, as in Ref. Fiaschi:2021okg the impact of the forward-backward $A_{\rm{FB}}$ asymmetry in the neutral current sector may be combined with that of the lepton-charge $A_{W}$ asymmetry in the charged current sector. This points to strategies which are alternative to those taken in experimental analyses such as in Refs. CMS:2018ktx ; ATLAS:2018gqq ; Bodek:2015ljm , and aim at exploiting new measurements, capable of providing sensitivity to PDFs with low theoretical and experimental systematics while controlling correlations. Related investigations of the $A_{\rm{FB}}$ asymmetry in Ref. Ball:2022qtp focus on the behavior induced by the NNPDF4.0 set NNPDF:2021njg . See also the studies Fu:2023rrs ; Yang:2022zvx ; Yang:2022bxv ; Xie:2022tzo ; Yang:2021cpd ; Fu:2020mxl based on the package ePump Schmidt:2018hvu ; Willis:2018yln .
@@ 4 Results
which follows the one used in Ref. H1:2015ubc . Here, $\mu^{i}$ is the measured value in bin $i$, $\delta_{i}$ is its experimental uncertainty, $m^{i}$ is the theoretical prediction, $\gamma^{i}_{j}$ is its relative uncertainty due to the PDF eigenvector $j$ which is shifted in units of sigma by $s_{j}$. This treatment of the PDF uncertainties follows the so-called profiling technique Paukkunen:2014zia ; HERAFitterdevelopersTeam:2015cre . In this method, the PDF uncertainties are included in the $\chi^{2}$ using nuisance parameters which are further constrained according to the tolerance criterion of the fit. As the tolerance criterion we use $\Delta\chi^{2}=1$. In this approach, one assumes that the new data are compatible with the theoretical predictions using the existing PDF set. No further theoretical uncertainties beyond the PDF uncertainties are considered when calculating the $\chi^{2}$. The uncertainties on the fitted parameters are obtained using the HESSE method which computes numerically the second derivatives of the $\chi^{2}$ with respect to the fitted parameters Alekhin:2014irh ; xfittergitlab . This assumes that the dependence of the theoretical predictions on the parameters of interest is linear near the minimum of the $\chi^{2}$. We cross-checked the uncertainties using the MINOS algorithm which uses the profile likelihood method to compute asymmetric confidence intervals, as well as the MNCONT algorithm James:1975dr which explicitly finds 2D contours where the $\chi^{2}$ is minimal, and found a good agreement with the hessian uncertainties.
@@
The results of the fit are presented in Figs. 1119 as allowed regions for different pairs of corrections to the $Z$ couplings to $u$- and $d$-type quarks at confidence level (CL) of 68%. Note that we fit four couplings at a time, while for presentation purposes we show 2D projection plots with different pairs of couplings. The resulting uncertainties on the couplings to the $d$-type quarks are roughly a factor of two larger than the corresponding uncertainties for the $u$-type quarks. A similar impact of the $A_{\text{FB}}$ measurements on the PDFs was found in Ref. Accomando:2019vqt : it was shown that these measurements are most sensitive to the weighted sum $(2/3)u_{v}+(1/3)d_{v}$ of the valence $u$- and $d$-quarks, and we report the same finding in the present study. This is related to the valence quark content of the proton and also to the fact that the $d\bar{d}$ initiated process gets more suppressed at high rapidity (where the sensitivity to the couplings is largest) in comparison to the $u\bar{u}$ initiated ones Accomando:2018nig . The exact details of this effect depend on the quark PDF behaviour at high values of the partonic momentum fraction $x$ Fiaschi:2022wgl ; Fu:2023rrs ; Hammou:2023heg ; Alekhin:2023uqx ; Ball:2022qtp .
@@
In Figs. 16 and 17 we compare our results obtained for the HL-LHC [^4] We show the results obtained using the ABMP16 PDF set, since only this set provides symmetric PDF errors which are easier to include in the PDF profiling. with the other analyses of existing data from LEP, Tevatron, HERA and LHC. The results are presented using either couplings to right- and left-handed quarks or axial and vector couplings. Namely, we compare with the analysis of the H1 Collaboration at HERA H1:2018mkk , the LEP+SLD combination ALEPH:2005ab , the analysis of D0 Collaboration at Tevatron D0:2011baz and the analysis of LEP, ATLAS and D0 data from Ref. Breso-Pla:2021qoe . [^5] Note that on these plots we compare published results, which are obtained with different methodology in some cases. E.g. in our work and Refs. Britzger:2020kgg ; Britzger:2022abi the assumption $\delta g_{R,L}^{Zd}=\delta g_{R,L}^{Zs}=\delta g_{R,L}^{Zb}$, $\delta g_{R,L}^{Zu}=\delta g_{R,L}^{Zc}$ was used, while in Ref. Breso-Pla:2021qoe the couplings to quarks of different flavours were fitted separately. Also, Refs. ALEPH:2005ab and Breso-Pla:2021qoe differ in the treatment of higher-order interference effects between the SM and new physics, universality assumption for down-type quarks, and the number of fitted parameters. In addition to the HL-LHC results, we present our results of analysing all available 10 bins from the ATLAS measurement of $A_{\text{FB}}$ ATLAS:2018gqq , while only $4$ bins at the $Z$ peak were used in the analysis of Ref. Breso-Pla:2021qoe . For the analysis of ATLAS data, we set the central data points to the theoretical predictions obtained at LO, while we use the data statistical and systematic uncertainties, as well as the PDF uncertainties. The analysis of ATLAS data follows the procedure used for the HL-LHC pseudodata. Such a procedure provides credible uncertainties on the $Z$ couplings, while in order to get meaningful central values one would need to use theoretical calculations with higher-order QCD corrections. Given that the experimental uncertainties of the ATLAS data are larger than the uncertainties of the HL-LHC pseudodata, the PDF uncertainties play a moderate role in this case. The level of precision expected at the HL-LHC outperforms any existing data sets ALEPH:2005ab ; Breso-Pla:2021qoe ; D0:2011baz ; H1:2018mkk ; ZEUS:2016vyd ; Abt:2016zth . Also, we note that using all the $10$ data bins from the ATLAS measurement ATLAS:2018gqq provides more information on some of the couplings (e.g. $\delta g_{R}^{Zd}$) compared to the $4$ bins at the $Z$ peak together with the LEP and D0 data which were used in the analysis of Ref. Breso-Pla:2021qoe .

### v3

@@ Abstract
Neutral current Drell-Yan (DY) lepton-pair production is considered to study $Z$-boson quark couplings. Using the open-source fit platform xFitter, we investigate the impact of high-statistics measurements of the neutral current DY (NCDY) forward-backward asymmetry $A_{\rm{FB}}$ near the weak boson mass scale in the present and forthcoming stages of the Large Hadron Collider (LHC). Besides recovering earlier results on the $A_{\rm{FB}}$ sensitivity to parton distribution functions, we analyze the precision determination of $Z$-boson couplings to left-handed and right-handed $u$-quarks and $d$-quarks, and explore Beyond-Standard-Model contributions using the Standard Model Effective Field Theory (SMEFT) framework. We perform a sensitivity study and comment on the role of the $A_{\rm{FB}}$ asymmetry for the electroweak SMEFT fit and precision $Z$-boson physics at the LHC and high-luminosity HL-LHC.
@@ 1 Introduction
Furthermore, as in Ref. Fiaschi:2021okg the impact of the forward-backward $A_{\rm{FB}}$ asymmetry in the neutral current sector may be combined with that of the lepton-charge $A_{W}$ asymmetry in the charged current sector. This points to strategies which are alternative to those taken in experimental analyses such as in Refs. CMS:2018ktx ; ATLAS:2018gqq ; Bodek:2015ljm , and aim at exploiting new measurements, capable of providing sensitivity to PDFs with low theoretical and experimental systematics while controlling correlations. Related investigations of the $A_{\rm{FB}}$ asymmetry in Ref. Ball:2022qtp focus on the behaviour induced by the NNPDF4.0 set NNPDF:2021njg . See also the studies Fu:2023rrs ; Yang:2022zvx ; Yang:2022bxv ; Xie:2022tzo ; Yang:2021cpd ; Fu:2020mxl based on the package ePump Schmidt:2018hvu ; Willis:2018yln .
@@ 4 Results
which follows the one used in Ref. H1:2015ubc . Here, $\mu^{i}$ is the measured value in bin $i$, $\delta_{i}$ is its experimental uncertainty, $m^{i}$ is the theoretical prediction, $\gamma^{i}_{j}$ is its relative uncertainty due to the PDF eigenvector $j$ which is shifted in units of sigma by $s_{j}$. This treatment of the PDF uncertainties follows the so-called profiling technique Paukkunen:2014zia ; HERAFitterdevelopersTeam:2015cre . In this method, the PDF uncertainties are included in the $\chi^{2}$ using nuisance parameters $s_{j}$ which are further constrained according to the tolerance criterion of the fit. The number of nuisance parameters $s_{j}$ corresponds to the number of eigenvectors for each PDF set. While $s_{j}$ are free parameters when minimizing the $\chi^{2}$, they do not change the number of degrees of freedom, because for each $s_{j}$ the corresponding prior is added, represented by the last term in Eq. (11), which acts as a data point. As the tolerance criterion we use $\Delta\chi^{2}=1$. In this approach, one assumes that the new data are compatible with the theoretical predictions using the existing PDF set. No further theoretical uncertainties beyond the PDF uncertainties are considered when calculating the $\chi^{2}$. The uncertainties on the fitted parameters are obtained using the HESSE method which computes numerically the second derivatives of the $\chi^{2}$ with respect to the fitted parameters Alekhin:2014irh ; xfittergitlab . This assumes that the dependence of the theoretical predictions on the parameters of interest is linear near the minimum of the $\chi^{2}$. We cross-checked the uncertainties using the MINOS algorithm which uses the profile likelihood method to compute asymmetric confidence intervals, as well as the MNCONT algorithm James:1975dr which explicitly finds 2D contours where the $\chi^{2}$ is minimal, and found a good agreement with the hessian uncertainties.
@@
The results of the fit are presented in Figs. 1119 as allowed regions for different pairs of corrections to the $Z$ couplings to $u$- and $d$-type quarks at confidence level (CL) of 68%.Note that we fit four couplings at a time, while for presentation purposes we show 2D projection plots with different pairs of couplings. The resulting uncertainties on the couplings to the $d$-type quarks are roughly a factor of two larger than the corresponding uncertainties for the $u$-type quarks. A similar impact of the $A_{\text{FB}}$ measurements on the PDFs was found in Ref. Accomando:2019vqt : it was shown that these measurements are most sensitive to the weighted sum $(2/3)u_{v}+(1/3)d_{v}$ of the valence $u$- and $d$-quarks, and we report the same finding in the present study. This is related to the valence quark content of the proton and also to the fact that the $d\bar{d}$ initiated process gets more suppressed at high rapidity (where the sensitivity to the couplings is largest) in comparison to the $u\bar{u}$ initiated ones Accomando:2018nig . The exact details of this effect depend on the quark PDF behaviour at high values of the partonic momentum fraction $x$ Fiaschi:2022wgl ; Fu:2023rrs ; Hammou:2023heg ; Alekhin:2023uqx ; Ball:2022qtp . We find a strong correlation (up to 0.95) between the different couplings, as illustrated in Figs. 1119, while the correlation coefficients between the couplings and the quark PDFs are moderate (about $0.5$). Furthermore, the latter exhibit significant variability across different values of $x$, reflecting the complex correlation of the different parton distributions in the original PDF sets.
@@
In Figs. 16 and 17 we compare our results obtained for the HL-LHC [^4] We show the results obtained using the ABMP16 PDF set, since only this set provides symmetric PDF errors which are easier to include in the PDF profiling. with the other analyses of existing data from LEP, Tevatron, HERA and LHC. The results are presented using either couplings to right- and left-handed quarks or axial and vector couplings. Namely, we compare with the analysis of the H1 Collaboration at HERA H1:2018mkk , the LEP+SLD combination ALEPH:2005ab , the analysis of D0 Collaboration at Tevatron D0:2011baz and the analysis of LEP, ATLAS and D0 data from Ref. Breso-Pla:2021qoe . [^5] Note that on these plots we compare published results, which are obtained with different methodology in some cases. E.g. in our work and Refs. Britzger:2020kgg ; Britzger:2022abi the assumption $\delta g_{R,L}^{Zd}=\delta g_{R,L}^{Zs}=\delta g_{R,L}^{Zb}$, $\delta g_{R,L}^{Zu}=\delta g_{R,L}^{Zc}$ was used, while in Ref. Breso-Pla:2021qoe the couplings to quarks of different flavours were fitted separately. Also, Refs. ALEPH:2005ab and Breso-Pla:2021qoe differ in the treatment of higher-order interference effects between the SM and new physics, universality assumption for down-type quarks, and the number of fitted parameters. In addition to the HL-LHC results, we present our results of analyzing all available 10 bins from the ATLAS measurement of $A_{\text{FB}}$ ATLAS:2018gqq , while only $4$ bins at the $Z$ peak were used in the analysis of Ref. Breso-Pla:2021qoe . For the analysis of ATLAS data, we set the central data points to the theoretical predictions obtained at LO, while we use the data statistical and systematic uncertainties, as well as the PDF uncertainties. The analysis of ATLAS data follows the procedure used for the HL-LHC pseudodata. Such a procedure provides credible uncertainties on the $Z$ couplings, while in order to get meaningful central values one would need to use theoretical calculations with higher-order QCD corrections. Given that the experimental uncertainties of the ATLAS data are larger than the uncertainties of the HL-LHC pseudodata, the PDF uncertainties play a moderate role in this case. The level of precision expected at the HL-LHC outperforms any existing data sets ALEPH:2005ab ; Breso-Pla:2021qoe ; D0:2011baz ; H1:2018mkk ; ZEUS:2016vyd ; Abt:2016zth . Also, we note that using all the $10$ data bins from the ATLAS measurement ATLAS:2018gqq provides more information on some of the couplings (e.g. $\delta g_{R}^{Zd}$) compared to the $4$ bins at the $Z$ peak together with the LEP and D0 data which were used in the analysis of Ref. Breso-Pla:2021qoe .